OpenStudy (sarahc):

Solve. 2x2 + 5x = 12. x=3/2 and x = –4 x=2 and x=-3 x = –6 and x = 5 x=1/2 and x=1/3

OpenStudy (anonymous):

2x^2 + 5x =12 Is this the equation?

OpenStudy (anonymous):

If so then it's x=3/2 and x = –4

OpenStudy (sarahc):

yes it is thank you!!

OpenStudy (anonymous):

ur welcome thks for the medal!